### v2

# On the Stability of Undesirable Equilibria in the Quadratic Program Framework for Safety-Critical Control

Abstract
Control Lyapunov functions (CLFs) and Control Barrier Functions (CBFs) have been used to develop provably safe controllers by means of quadratic programs (QPs). This framework guarantees safety in the form of trajectory invariance with respect to a given set, but it can introduce undesirable equilibrium points to the closed loop system, which can be asymptotically stable. In this work, we present a detailed study of the formation and stability of equilibrium points with the CLF-CBF-QP framework with multiple CBFs. In particular, we prove that undesirable equilibrium points occur for most systems, and their stability is dependent on the CLF and CBF geometrical properties. We introduce the concept of CLF-CBF compatibility for a system, regarding a CLF-CBF pair inducing no stable equilibrium points other than the CLF global minimum on the corresponding closed-loop dynamics. Sufficient conditions for CLF-CBF compatibility for LTI and drift-less full-rank systems with quadratic CLF and CBFs are derived, and we propose a novel control strategy to induce smooth changes in the CLF geometry at certain regions of the state space in order to satisfy the CLF-CBF compatibility conditions, aiming to achieve safety with respect to multiple safety objectives and quasi-global convergence of the trajectories towards the CLF minimum. Numeric simulations illustrate the applicability of the proposed method.
Index Terms:
Lyapunov methods, Control barrier functions
I Introduction
The engineering of safety-critical systems is a fruitful and rich topic receiving a growing amount of attention nowadays. Safety-critical systems are of crucial importance for many industrial sectors and production lines, where the stability of feedback-controlled systems is just so important as their capacity to provide safe behaviour under a wide variety of operational circumstances. Furthermore, safety is also a mandatory property for systems with high levels of interoperability, cooperation, or coordination with humans.
The notion of safety was first introduced in 1977 in the context of program correctness by [1] and later formalized in [2], which also introduced the concept of liveness. Intuitively, one can describe these two contrasting system properties as: (i) the requirement of avoiding undesired situations while (ii) guaranteeing the eventual achievement of a desired configuration, respectively. As pointed out by [3], in the context of control systems, liveness can be identified as an asymptotic stability requirement with respect to a certain set of desired or objective states, while safety can be defined as the invariance of the system trajectories to some set, defined as the set of safe states. While the design of asymptotically stabilizing controllers has been extensively studied in control Lyapunov theory [4], the design of controllers capable of guaranteeing safety has been the subject of study in the topic of Control Barrier Functions (CBFs) [5]. Additionally, [5] also introduced the idea of unifying CBFs with Control Lyapunov Functions (CLFs) through the use of quadratic programs (QPs), combining safety and stabilization requirements in a single control framework.
However, the study of controllers combining the two desirable properties of stability and safety is still in early stages. In [6], it is shown that the QP-based framework proposed by [5] can introduce undesirable equilibrium points other than the CLF minimum into the closed-loop system. The fact that some of these undesirable equilibrium points can be asymptotically stable and can be arbitrarily close to the set of unsafe states is an important practical limitation of the framework, since it could result in system deadlocks and expose the system to close-to-failure situations, forcing the designer to opt for highly conservative safety margins when designing the system safety specifications. In [7], a CBF-based controller was proposed in which safety is ensured with respect to multiple non-convex unsafe regions and undesirable stable equilibrium are practically avoided, but the method is dependent on the computation of a nonlinear “convexification” function for the unsafe sets, which is dependent on the barrier geometry and could be computationally hard to solve. Furthermore, it is not clear how these results can be generalized for the CLF-CBF framework. In [8], the problem of deadlocks in the QP-based formulation for safety-critical systems was addressed for the safety-filter CBF-QP-based controller as proposed in [3]. In this context, deadlocks are caused by a conflict between the stabilization objectives of the nominal controller and the safety barriers, and were managed by introducing a consistent perturbation into the QP constraints. Although efficient for solving some types of deadlocks, this proposed method modifies the safety constraints and allows for the possibility of leaving the safe set if the deadlock situation happens to occur on the boundary, which can lead to unsafe behavior. Considering the CLF-CBF framework, [9] has proposed a modified CLF-CBF-based QP controller in which interior equilibrium points and certain boundary equilibrium points satisfying a certain condition do not exist for the resulting closed-loop system. However, boundary equilibrium points could still occur in general.
The contribution of the present work are as follows:
  1. i
    We present an analysis for the conditions for existence and stability of all types of equilibrium points occurring in CLF-CBF-QP-based framework [5], for the class of nonlinear control affine systems.
  2. ii
    We introduce the concept of CLF-compatibility, denoting the property of a CLF that, for a given system dynamics and fixed set of CBFs modeling the safety requirements for the task, the CLF-CBF-QP controller does not introduce any stable equilibrium points other than the CLF global minimum.
  3. iii
    We derive necessary and sufficient conditions for CLF compatibility for the classes of (i) control linear systems with full-rank input matrix and (ii) linear time-invariant (LTI) systems, with quadratic CLF and CBFs.
  4. iv
    We propose a method for computing a corresponding compatible CLF from a non-compatible one, and a CLF-CBF-QP controller that adaptively modifies the CLF geometry to achieve CLF-compatibility using the found compatible CLF, thereby guaranteeing safety and quasi-global convergence of the closed-loop system trajectories.
II Preliminary
Notation: The fields of real and complex numbers are $\mathbb{R}$ and $\mathbb{C}$, respectively. Given a matrix $A\in\mathbb{R}^{n\times m}$, $[A]_{ij}\in\mathbb{R}$ denotes its $i$-th row, $j$-th column component and $[A]_{k}\in\mathbb{R}^{n}$ denotes its $k$-th column. The group of real symmetric matrices is $\mathcal{S}^{n}\subset\mathbb{R}^{n\times n}$. The determinant of a square matrix $A$ is $|A|$, its Frobenius norm is $\lVert A\rVert_{\mathcal{F}}$, and its adjoint matrix is $\operatorname{adj}{A}\in\mathbb{R}^{n\times n}$, where $A\,\operatorname{adj}{A}=|A|I_{n}$, where $I_{n}\in\mathbb{R}^{n\times n}$ is the $n\times n$ identity matrix. Given a vector $v\in\mathbb{R}^{n}$, $[v]_{k}\in\mathbb{R}$ is its $k$-th component. A scalar-valued function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is said to be of (differentiability) class $C^{k}$ if all of its $k$-th order partial derivatives exist and are continuous. Consider the class $C^{2}$ function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$: (i) its gradient is defined as the vector-valued function $\nabla f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ such that $[\nabla f(x)]_{k}=\frac{\partial f(x)}{\partial x_{k}}=\partial_{k}f(x)$, where $\partial_{k}$ denotes partial differentiation with respect to the $k$-th component of the function input, (ii) its Hessian matrix is defined as the matrix-valued function $H_{f}:\mathbb{R}^{n}\rightarrow\mathcal{S}^{n}$ such that $[H_{f}(x)]_{ij}=\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}$. $L_{g}f$ is the Lie derivative of $f$ along $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$, that is, $L_{g}f=\nabla f^{\mathsf{T}}g\in\mathbb{R}^{m}$. The inner product between two vectors $u,v\in\mathbb{R}^{n}$ induced by a positive semidefinite matrix $G=G^{\mathsf{T}}\geq 0$ is given by $\langle u,v\rangle_{G}=u^{\mathsf{T}}G\,v$. This inner product induces a norm $\lVert v\rVert_{G}^{2}=\langle v,v\rangle_{G}=v^{\mathsf{T}}G\,v$ over $\mathbb{R}^{n}$. The standard inner product is then $\langle u,v\rangle=\langle u,v\rangle_{I_{n}}$, with standard Euclidean norm $\lVert v\rVert^{2}=\lVert v\rVert_{I_{n}}^{2}$. The orthogonal complement of a subspace $\mathcal{W}$ is denoted by $\mathcal{W}^{\perp}$, with the notion of orthogonality dependent upon the considered inner product $\langle\cdot,\cdot\rangle_{G}$. The set $\text{span}\{v_{1},\cdots,v_{p}\}$ is the set of all linear combinations of vectors from $\{v_{1},\cdots,v_{p}\}\subset\mathbb{R}^{n}$. The positive semi-definite cone of symmetric matrices is $\mathcal{S}^{n}_{+}$. The null space and spectrum of a real square matrix $A\in\mathbb{R}^{n\times n}$ are given by $\mathcal{N}(A)\subset\mathbb{R}^{n}$ and $\sigma(A)\subset\mathbb{R}$, respectively.
Consider the nonlinear control affine system
$\dot{x}=f(x)+g(x)u$ (1)
where $x\in\mathbb{R}^{n}$ is the system state and $u\in\mathbb{R}^{m}$ is the control input. Vector fields $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ are locally Lipschitz.
Definition II.1 (CLFs).
A positive definite function $V$ is a control Lyapunov function (CLF) for system (1) if it satisfies:
$\inf_{u\in\mathbb{R}^{m}}\left[L_{f}V(x)+\langle L_{g}V(x),u\rangle\right]\leq-\gamma(V(x))$
where $\gamma:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is a class $\mathcal{K}$ function [4].
This definition implies that there exists a set of stabilizing controls that makes the CLF strictly decreasing everywhere outside its global minimum $x_{0}\in\mathbb{R}^{n}$.
Definition II.2 (Safety).
The trajectories of a given system are safe with respect to a set $\mathcal{C}$ if $\mathcal{C}$ is forward invariant, meaning that for every $x(0)\in\mathcal{C}$, $x(t)\in\mathcal{C}$ for all $t>0$.
Consider $N$ subsets $\mathcal{C}_{1},\dots,\mathcal{C}_{N}\subset\mathbb{R}^{n}$ defined by the superlevel set of a continuously differentiable function $h_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$:
$\mathcal{C}_{i}$ =$=\{x\in\mathbb{R}^{n}:h_{i}(x)\geq 0\},\quad i=1,2,\ldots,N$ (2)
Definition II.3 (CBFs).
Let $\mathcal{C}_{i}\in\mathbb{R}^{n}$ be defined by one of the functions (2). Then $h_{i}(x)$ is a (zeroing) Control Barrier Function (CBF) for (1) if there exists a locally Lipschitz extended class $\mathcal{K}_{\infty}$ function [^1] An extended class $\mathcal{K}_{\infty}$ function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ is strictly increasing with $\alpha(0)\!=\!0$. $\alpha$ such that
$\sup_{u\in\mathbb{R}^{m}}\left[L_{f}h_{i}(x)+\langle L_{g}h_{i}(x),u\rangle\right]\geq-\alpha(h_{i}(x))\quad\forall x\in\mathbb{R}^{n}\,.$
This definition simply means that the CBFs $h_{i}(x)$ are only allowed to decrease in the interior of their respective safe sets $\text{int}(\mathcal{C}_{i})$, but not on their boundaries $\partial\mathcal{C}_{i}$.
Consider the closed-loop system for (1)
$\dot{x}=f_{cl}(x):=f(x)+g(x)u^{\star}(x)$ (3)
with control law $u^{\star}(x)$ given by the minimum-norm feedback controller based on [5]
$u^{\star}(x)$ $=\operatorname*{argmin}_{(u,\delta)\in\mathbb{R}^{m+1}}\frac{1}{2}\lVert u\rVert^{2}+\frac{1}{2}p\delta^{2}\qquad$ (4)
$s.t.\,L_{f}V$ $+\langle L_{g}V,u\rangle+\gamma(V)\leq\delta$
$\quad L_{f}h_{i}$ $+\langle L_{g}h_{i},u\rangle\geq-\alpha(h_{i})\,,\quad i\in\{1,\cdots,N\}$
with $p>0$, $\gamma$ and $\alpha$ being class $\mathcal{K}$ and class $\mathcal{K}_{\infty}$ functions, respectively. If feasible, the feedback controller (4) guarantees local stability of $x_{0}$ and safety of the closed-loop system trajectories with respect to the safe set
$\mathcal{C}=\cap^{N}_{i=1}\mathcal{C}_{i}$ (5)
However, (4) does not guarantee global stabilization, meaning that the trajectories could converge towards equilibrium points other than the CLF minimum [6].
Assumption II.1.
The initial state $x(0)\in\mathbb{R}^{n}$ is contained in the safe set (5) and the CLF minimum $x_{0}$ is contained in $\mathcal{C}$, that is, $h_{i}(x_{0})\geq 0$ for all $i\in\{1,\cdots,N\}$.
Assumption II.1 comes from the fact that it is natural to assume that a system starts in a safe configuration; as an example, it is only natural to assume that a vehicle starts its navigation task in a safe state of non-collision against obstacles. Furthermore, the CLF minimum $x_{0}$ must be reachable by the controller (4).
Theorem 1.
Under Assump. II.1, the QP (4) is feasible for all $x\in\mathcal{C}$, if at least one the two conditions are met:
  • (i)
    There is only one CBF ($N=1$).
  • (ii)
    The affine non-linear system (1) is driftless, that is, $f(x)=0\,\,\,\forall x\in\mathbb{R}^{n}$.
Proof.
The proof for (i) was first introduced in [5]. The proof of (ii) is as follows: under Assumption II.1, the initial state $x(0)\in\mathcal{C}$, that is, $h_{i}(x(0))\geq 0$ for all $i=1,\dots,N$. Then, for driftless affine nonlinear systems, the decision space associated to the $i$-th CBF constraint is given by the half-plane $\mathbb{K}_{cbf_{i}}(x)=\{(u,\delta)\in\mathbb{R}^{m+1}:\langle L_{g}h_{i}(x),u\rangle+\alpha h_{i}(x)\geq 0\}$. The intersection of these half-planes configures a convex polytope $\mathbb{K}_{cbf}(x)=\cap^{N}_{i=1}\mathbb{K}_{cbf_{i}}(x)$, which is the decision space associated to the CBF constraints. Due to the independence of the CBF constraints on the slack variable $\delta$ and due to the fact that $h_{i}(x(0))\geq 0$, $\mathbb{K}_{cbf}(x(0))$ contains the the entire $\delta$-axis, that is, the line $(u,\delta)=(0,\delta)\,\,\forall\delta\in\mathbb{R}$. Therefore, $\mathbb{K}_{cbf}(x(0))$ is unbounded and non-empty. Defining the decision space associated to the CLF constraint as the half-plane $\mathbb{K}_{clf}(x)=\{(u,\delta)\in\mathbb{R}^{m+1}:\langle L_{g}V(x),u\rangle+\gamma(V(x))\leq\delta\}$, the feasible set associated to the QP (4) is the intersection $\mathbb{K}_{clf}(x)\cap\mathbb{K}_{cbf}(x)\subset\mathbb{R}^{m+1}$. Notice that $\mathbb{K}_{clf}(x(0))\cap\mathbb{K}_{cbf}(x(0))\neq\emptyset$, meaning that the QP is initially feasible under Assumption II.1. Then, since the CBF constraints guarantee the invariance of the trajectories $x(t)$ with respect to the safe set $\mathcal{C}$, $h_{i}(x(t))\geq 0\forall t\geq 0$, $i=1,\cdots,N$. Therefore, the convex polytope $\mathbb{K}_{cbf}(x(t))$ remains unbounded and non-empty $\forall t\geq 0$ (it must always contain the $\delta$-axis), and therefore the feasible set for the QP (4) is $\mathbb{K}_{clf}(x(t))\cap\mathbb{K}_{cbf}(x(t))\neq\emptyset$ for all $t\geq 0$. $\square$
Assumption II.2 (Disjoint Unsafe Sets).
The unsafe sets of the $N$ barriers are disjoint, that is:
$\overline{\mathcal{C}}_{i}\cap\overline{\mathcal{C}}_{j}=\emptyset\qquad\forall i\neq j$ (6)
Remark II.1.
Assumption II.2 is not restrictive for the following reason: assume there exist barriers $h_{i}$, $h_{j}$ with non-empty unsafe set intersections, that is, $\overline{\mathcal{C}}_{i}\cap\overline{\mathcal{C}}_{j}\neq\emptyset$. Then, it is possible to construct a new composite barrier $h_{k}$ with unsafe set $\overline{\mathcal{C}}_{k}\supset\overline{\mathcal{C}}_{i}\cup\overline{\mathcal{C}}_{j}$ (under mild assumptions on the regularities of $h_{k}$), thus representing (almost) the same safe region as $\mathcal{C}_{i}\cap\mathcal{C}_{j}$. [10] proposes a composition method for combining multiple CBFs into a single one.
Definition II.4.
Given a CLF $V:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$, define the transformed CLF $\overline{V}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ as
[…]
$\nabla\overline{V}$ $=\gamma(V)\nabla V$ (8)
$H_{\overline{V}}$ $=\gamma(V)H_{V}+\gamma^{\prime}(V)\nabla V\nabla V^{\mathsf{T}}$ (9)
Proposition 1.
The transformed CLF (7) has the following properties:
  • (i)
    $\overline{V}(x)>0$ $\forall x\neq x_{0}$. Additionally, $\overline{V}(x_{0})=0$.
  • (ii)
    The integral transformation of (7) is invertible.
  • (iii)
    $V(x)$ and $\overline{V}(x)$ have the same level sets.
Proof.
Property (i) can be seen from the fact that $\gamma:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is a class $\mathcal{K}$ function, and therefore its integral is positive and strictly increasing. Furthermore, at $x=x_{0}$, $V(x_{0})=0$ and the limits of integration on (7) are both zero, showing that $\overline{V}(x_{0})=0$. Property (ii) can also be inferred from the fact that $\gamma$ is of class $\mathcal{K}$: since its integral is a positive and strictly increasing function, its inverse always exists. That means that the original $V$ can always be computed from the transformed CLF $\overline{V}$ by inverting the integral transformation (7). Property (iii) holds because by (8), the gradients of $V$ and $\overline{V}$ are co-directed and $V$, $\overline{V}$ are continuous functions. Therefore, they must share the same level sets. $\square$
As will be shown in the next sections, the CLF transformation in Definition II.4 will be useful not only for expressing the existence and stability conditions for equilibrium points in a simpler way, but also for developing the method for CLF-compatibility that is presented in Section V.
III Existence of Equilibrium Points in the CLF-CBF Framework
In this section, we extend a result from [6], regarding the existence of equilibrium points when multiple CBF constraints are present.
Definition III.1 (Equilibrium Manifold).
Define the vector-valued transformation $f_{i}:\mathbb{R}^{n}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n\times n}$ associated to the $i$-th CBF as
$f_{i}(x,\lambda)=f+\lambda G\nabla h_{i}-pG\nabla\overline{V}$ (10)
The Jacobian matrix of $f_{i}$ with respect to $x\in\mathbb{R}^{n}$ is
$J_{f_{i}}(x,\lambda)=$ $=\frac{\partial f}{\partial x}+\lambda\frac{\partial G\nabla h_{i}}{\partial x}-p\frac{\partial G\nabla\overline{V}}{\partial x}$ (11)
As will be demonstrated in the next sections, (10)-(11) will be of central importance to characterize the existence and stability conditions for the equilibrium points of the closed-loop system.
Theorem 2 (Existence of Equilibrium Points).
Let (3) be the closed-loop system formed by the nonlinear system (1) with Assumption II.2 and controller (4). The set $\mathcal{E}$ of equilibrium points of (3) is given by
$\mathcal{E}$ $=\left(\bigcup^{N}_{i=1}\mathcal{E}_{\partial\mathcal{C}_{i}}\right)\cup\mathcal{E}_{int(\mathcal{C})}$ (12)
where $\mathcal{E}_{\partial\mathcal{C}_{i}}=\partial\mathcal{C}_{i}\cap\{x\in\mathbb{R}^{n}\,|\,\exists\lambda\geq 0\text{ s.t. }f_{i}(x,\lambda)=0\}$ is the set of boundary equilibrium points and $\mathcal{E}_{int(\mathcal{C})}=int(\mathcal{C})\cap\{x\in\mathbb{R}^{n}\,|\,f(x)=pG(x)\nabla\overline{V}(x)\}$ is the set of interior equilibrium points.
Proof.
The Lagrangian associated to QP (4) is
$\mathcal{L}=0.5\left(\lVert u\rVert^{2}+p\delta^{2}\right)+\lambda_{0}$ $(F_{V}+L_{g}Vu-\delta)$
$-\sum^{N}_{i=1}$ $\lambda_{i}(F_{h_{i}}+L_{g}h_{i}\,u)$ (13)
where $F_{V}(x)=L_{f}V+\gamma(V)$ and $F_{h_{i}}(x)=L_{f}h_{i}+\alpha(h_{i})$, and $\lambda_{i}\geq 0\in\{0,1,\cdots,N\}$ are the KKT multipliers associated to the optimization problem. Then, the KKT conditions are:
$\frac{\partial\mathcal{L}}{\partial u}=u+\lambda_{0}g^{\mathsf{T}}\nabla V-\sum^{N}_{i=1}\lambda_{i}g^{\mathsf{T}}\nabla h_{i}$ $=0$ (14)
$\frac{\partial\mathcal{L}}{\partial\delta}=p\delta-\lambda_{0}$ $=0$ (15)
$\lambda_{0}(F_{V}+L_{g}V\,u-\delta)$ $=0$ (16)
$\lambda_{i}(F_{h_{i}}+L_{g}h_{i}\,u)$ $=0$ (17)
with $i\in\{1,\cdots,N\}$. Using (14)-(15), the QP solutions are given by:
$u^{\star}(x)$ $=g^{\mathsf{T}}\left(-\lambda_{0}\nabla V+\sum^{N}_{i=1}\lambda_{i}\nabla h_{i}\right)$ (18)
$\delta^{\star}(x)$ $=p^{-1}\lambda_{0}\,,$ (19)
with $\lambda_{i}\geq 0\in\{0,1,\cdots,N\}$. Substituting (18) on (3) yields the following expression for the closed-loop system:
$f_{cl}(x)\!=\!f\!+\!G\left(-\lambda_{0}\nabla V\!+\!\sum^{N}_{i=1}\lambda_{i}\nabla h_{i}\right)\,,$ (20)
where $G(x)=g(x)g(x)^{\mathsf{T}}\in\mathbb{R}^{n\times n}\geq 0$ is a positive semi-definite matrix. At an equilibrium point $x_{e}\in\mathcal{E}$, $f_{cl}(x_{e})=0$. Applying this condition to (20) yields
$f(x_{e})=G(x_{e})\left(\lambda_{0}\nabla V(x_{e})-\sum^{N}_{i=1}\lambda_{i}\nabla h_{i}(x_{e})\right)$ (21)
with $\lambda_{i}\geq 0$.
Case 1. Consider the region of the state space where the CLF constraint is inactive: $L_{f_{cl}(x)}V(x)+\gamma(V(x))-\delta^{\star}(x)<0$. From (16), $\lambda_{0}=0$. Then, using (19), notice that $\delta^{\star}(x)=0$. At an equilibrium point $x_{e}\in\mathcal{E}$, $L_{f_{cl}(x_{e})}V(x_{e})=0$, and therefore we obtain $\gamma(V(x_{e}))<\delta^{\star}(x_{e})=0$, implying that $V(x_{e})<0$, which is a contradiction since $V$ is a nonnegative function. Therefore, all equilibrium points must lie on the region where the CLF constraint is active.
Case 2. Consider the region where CLF constraint is active: $L_{f_{cl}(x)}V(x)+\gamma(V(x))=\delta^{\star}(x)$. At an equilibrium point $x_{e}\in\mathcal{E}$, $L_{f_{cl}(x_{e})}V(x_{e})=0$. Therefore, using (19), $\gamma(V(x_{e}))=\delta^{\star}(x_{e})=p^{-1}\lambda_{0}$. Then, at any equilibrium point $x_{e}\in\mathcal{E}$, the KKT multiplier associated to the CLF constraint is $\lambda_{0}(x_{e})=p\gamma(V(x_{e}))\geq 0$. Therefore, equation (21) yields:
$f(x_{e})\!=\!G(x_{e})\!\left(\!p\nabla\overline{V}(x_{e})\!-\!\sum^{N}_{i=1}\lambda_{i}\nabla h_{i}(x_{e})\!\right)$ (22)
where $\nabla\overline{V}(x_{e})=\gamma(V(x_{e}))\nabla V(x_{e})$. For the next two cases, the CLF constraint is assumed to be active.
Case 3. Consider the region where the $i$-th CBF constraint is active: $L_{f_{cl}(x)}h_{i}(x)+\alpha(h_{i}(x))=0$. At $x_{e}\in\mathcal{E}$, $L_{f_{cl}(x_{e})}h_{i}(x_{e})=0$, implying that $h_{i}(x_{e})=0$. Therefore, equilibrium points occurring in this region must lie on the boundary of the $i$-th safe set, that is, $x_{e}\in\partial\mathcal{C}_{i}$. Next, we show that, under Assumption II.2, these equilibrium points can only occur when only the $i$-th CBF constraint is active. Assume that $x_{e}\in\mathcal{E}$ occurs when two CBF constraints are active: that is, we have $L_{f_{cl}(x_{e})}h_{i}(x_{e})+\alpha(h_{i}(x_{e}))=0$ and $L_{f_{cl}(x_{e})}h_{j}(x_{e})+\alpha(h_{j}(x_{e}))=0$, for $i\neq j$, and therefore, since $L_{f_{cl}(x_{e})}h_{i}(x_{e})=0$ and $L_{f_{cl}(x_{e})}h_{j}(x_{e})=0$, we have $h_{i}(x_{e})=h_{j}(x_{e})=0$. However, by Assumption II.2, this is a contradiction since $\partial\mathcal{C}_{i}\cap\partial\mathcal{C}_{j}=\emptyset$. The conclusion is that boundary equilibrium points on the $i$-th boundary are on the set where only the $i$-th CBF constraint is active, denoted by $\mathcal{S}_{i}$. That means that at $x_{e}\in\partial\mathcal{C}_{i}$, $\lambda_{j}=0\,,\,\,\forall j\neq i$. Therefore, (22) reduces to
$f(x_{e})=G(x_{e})\left(p\nabla\overline{V}(x_{e})-\lambda_{i}\nabla h_{i}(x_{e})\right)$ (23)
where $\lambda_{i}\geq 0$ is the corresponding KKT multiplier. Notice that (23) is equivalent to $f_{i}(x,\lambda_{i})=0$, with $f_{i}$ defined by (10). Thus, in this case, the equilibrium point is on the boundary of the safe set and satisfies $f_{i}(x_{e},\lambda_{i})$ for some $\lambda_{i}\geq 0$, proving the construction of $\mathcal{E}_{\partial\mathcal{C}_{i}}$.
Case 4. Consider the region where all CBF constraints are inactive: $L_{f_{cl}(x)}h_{i}(x)+\alpha(h_{i}(x))>0$. From (17), $\lambda_{1}=\cdots=\lambda_{N}=0$. At an equilibrium point $x_{e}\in\mathcal{E}$, $L_{f_{cl}(x_{e})}h_{i}(x_{e})=0$, implying that $h_{i}(x_{e})>0$. Therefore, equilibrium points occurring in this region must lie on the interior of the safe set, that is, $x_{e}\in int(\mathcal{C})$. Additionally, (22) must be satisfied with $\lambda_{1}=\cdots=\lambda_{N}=0$, which means that $f(x_{e})=pG(x_{e})\nabla\overline{V}(x_{e})$. This proves the construction of $\mathcal{E}_{int(\mathcal{C})}$. $\square$
A similar version of Theorem 2 was demonstrated in [6], considering only one CBF. Therefore, combining stabilization and safety objectives with the CLF-CBF framework can introduce equilibrium points in the closed-loop system other than the CLF global minimum $x_{0}\in\mathbb{R}^{n}$, some of them could even possibly be asymptotically stable [6]. This is a known problem in CLF-CBF literature and was considered in other works as well, such as in [9], which has presented a similar characterization of the equilibrium points and has proposed a modified QP-based controller for (1): $u(x)=\overline{u}(x)+u_{\text{nom}}(x)$, where $\overline{u}(x),u_{\text{nom}}(x)\in\mathbb{R}^{m}$ are feedback controllers to be designed as follows. Substituting $u(x)$ into (1) transforms the system dynamics into:
$\dot{x}$ $=\overline{f}(x)+g(x)\overline{u}(x)$ (24)
where $\overline{f}(x)=f(x)+g(x)u_{\text{nom}}(x)$.
Assumption III.1 (CLF Condition).
Given a nonlinear control affine dynamical system (1), the CLF $V(x)$ satisfies $L_{f}V(x)<0\,,\forall x\neq x_{0}\in\mathbb{R}^{n}$.
Let $u_{\text{nom}}(x)$ be a feedback controller chosen in such a way that Assump. III.1 holds for the CLF $V$ and system (24). This can always be done if the system (1) is controllable: (i) in case $V$ already satisfies Assump. III.1 for the original system (1), $u_{\text{nom}}(x)=0$ and $\overline{f}(x)=f(x)$, (ii) in case $V$ does not satisfy Assump. III.1 for the original system (1), assuming that the system is controllable, one can always find $u_{\text{nom}}(x)$ such that $V$ satisfies Assump. III.1 for $\overline{f}(x)=f(x)+g(x)u_{\text{nom}}(x)$ of the transformed system (24), for example, by using Sontag’s formula [11]. In [9], it was shown that with $u_{\text{nom}}(x)$ chosen in this way and with $\overline{u}(x)$ obtained from solving the QP (4) with the system model given by the transformed dynamics (24), then the closed-loop system obtained from applying controller $u(x)=\overline{u}(x)+u_{\text{nom}}(x)$ into (1) has the following set of equilibrium points:
$\overline{\mathcal{E}}$ $=\left(\bigcup^{N}_{i=1}\overline{\mathcal{E}}_{\partial\mathcal{C}_{i}}\right)\cup\{x_{0}\}$ (25)
¯$\overline{\mathcal{E}}_{\partial\mathcal{C}_{i}}$ $=\mathcal{E}_{\partial\mathcal{C}_{i}}\cap\{x\in\mathbb{R}^{n}\,|\,L_{g}h_{i}(x)\neq 0\}$
That is, interior equilibrium points other than the CLF minimum and boundary equilibrium points satisfying $L_{g}V(x)=0$ do not exist. However, the existence of boundary equilibrium points with $L_{g}V(x)\neq 0$ is not excluded. Since $\overline{u}(x)$ is obtained through solving the QP (4) for the a new nonlinear control affine system (24), the theory developed so far for the remaining closed-loop equilibrium points remains valid.
Proposition 2.
If Assump. III.1 holds for a CLF $V(x)$, it also holds for the transformed CLF $\overline{V}$ from Definition II.4.
Proof.
If Assump. III.1 holds for $V(x)$, then $L_{f}V=\langle\nabla V,f\rangle<0\,\forall x\neq x_{0}$. Using (8), for all $x\neq x_{0}$ we have $\nabla V=\gamma(V)^{-1}\nabla\overline{V}$ with $\gamma(V)^{-1}>0$. Then, $L_{f}V=\gamma(V)^{-1}\langle\nabla\overline{V},f\rangle<0$ implies $L_{f}\overline{V}<0$ for all $x\neq x_{0}$. $\square$
IV Stability of Equilibrium Points in the CLF-CBF Framework
Assuming system controllability, since it is always possible to apply the technique proposed in [9] and work with the transformed system (24), from here on we assume to be working with a nonlinear control affine system such that Assump. III.1 is satisfied. This way, only the remaining boundary equilibrium points with $L_{g}h(x)\neq 0$ need to be addressed. Our objective in this section is to study the stability properties of these points. Particularly, we generalize [6, Theorem 2] for nonlinear control affine systems, deriving a necessary and sufficient condition for the instability of boundary equilibrium points satisfying $L_{g}h(x)\neq 0$, when multiple CBF constraints are present.
Definition IV.1.
Define the two vector fields $z_{1},z_{2}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ associated to the $i$-th barrier as
$z_{1}(x)$ $=\frac{\nabla h_{i}}{\lVert\nabla h_{i}\rVert_{G}}\,,$ (26)
$z_{2}(x)$ $=\nabla V-\langle\nabla V,z_{1}\rangle_{G}\,z_{1}\,,$ (27)
where $G(x)=g(x)g(x)^{\mathsf{T}}$.
One can verify that $\{z_{1},z_{2}\}$ is an orthogonal set of vectors with respect to the inner product $\langle\cdot,\cdot\rangle_{G}$, that is, $\langle z_{i},z_{j}\rangle_{G}=0$, $i,j\in\{1,2\}$, $i\neq j$. In particular, $\langle z_{1},z_{1}\rangle_{G}=1$. Furthermore, define the scalar function $\eta:\mathbb{R}^{n}\rightarrow\mathbb{R}_{+}$ as
$\eta(x)$ $=(1+p\langle z_{2},z_{2}\rangle_{G})^{-1}$ (28)
with the following properties:
  1. (i)
    $0<\eta(x)\leq 1\,\,\forall x\in\mathbb{R}^{n}$, since $\langle z_{2},z_{2}\rangle_{G}\geq 0$.
  2. (ii)
    $\eta=1$ if and only if (i) $z_{2}=0$ or (ii) $Gz_{2}=0$.
  3. (iii)
    From (28), the inner product $\langle z_{2},z_{2}\rangle_{G}$ can be expressed as $\langle z_{2},z_{2}\rangle_{G}=p^{-1}(\eta^{-1}-1)\geq 0$,
  4. (iv)
    Combining (27) and (iii), it is possible to demonstrate $p^{-1}+\lVert\nabla V\rVert_{G}^{2}=\langle\nabla V,\hat{z}_{1}\rangle_{G}^{2}+p^{-1}\eta^{-1}$.
Lemma 1 (Boundary Jacobian).
Under Assumption II.2, the Jacobian matrix $J_{cl}(x_{e})\in\mathbb{R}^{n\times n}$ of the closed-loop system (3) computed at a boundary equilibrium point $x_{e}\in\mathcal{E}_{\partial\mathcal{C}_{i}}$ with $L_{g}h_{i}(x_{e})\neq 0$ and corresponding KKT multiplier $\lambda_{i}\geq 0$ is given by
$J_{cl}(x_{e})$ $\!=\!\left(I_{n}\!-\!GZN_{1}Z^{\mathsf{T}}\right)J_{i}(x_{e},\lambda_{i})\!-\!GZN_{1}\Psi Z^{\mathsf{T}}$ (29)
where $Z(x)=[\,z_{1}\,\,\,z_{2}\,]\in\mathbb{R}^{n\times 2}$ with $z_{i}(x)$ as defined in (26)-(27), $N_{1}(x)=\text{diag}\{1,p\eta(x)\}>0$ with $\eta(x)$ as defined in (28). Matrices $J_{i}$ and $\Psi$ are given by
$J_{i}(x,\lambda)$ $=\frac{\partial f(x)}{\partial x}+\lambda_{i}\frac{\partial G\nabla h_{i}}{\partial x}-p\gamma(V)\frac{\partial G\nabla V}{\partial x}$ (30)
$\Psi$ $=\begin{bmatrix}\alpha^{\prime}(h_{i})&0\\ \langle\nabla V,\hat{z}_{1}\rangle_{G}(\gamma^{\prime}(V)-\alpha^{\prime}(h_{i}))&\gamma^{\prime}(V)\end{bmatrix}$ (31)
Proof.
This demonstration is a direct continuation of the proof of Theorem 2. In Case 3, one can substitute (18) and (19) into the complementary slackness conditions (16)-(17) and use the fact that $\lambda_{j}=0$ for all $j\neq i$ to get the following system:
$\underbrace{\begin{bmatrix}p^{-1}+\lVert\nabla V\rVert_{G}^{2}\!&\!-\langle\nabla V,\nabla h_{i}\rangle_{G}\\ -\langle\nabla V,\nabla h_{i}\rangle_{G}\!&\!\lVert\nabla h_{i}\rVert_{G}^{2}\end{bmatrix}}_{C(x)}\underbrace{\begin{bmatrix}\lambda_{0}\\ \lambda_{i}\end{bmatrix}}_{\bar{\lambda}_{i}}\!=\!\underbrace{\begin{bmatrix}F_{V}\\ -F_{h_{i}}\end{bmatrix}}_{b(x)}$ (32)
where $F_{V}=L_{f}V+\gamma(V)$ and $F_{h_{i}}=L_{f}h_{i}+\alpha(h_{i})$. In particular, the set $\mathcal{S}_{i}$ where the CLF and only the $i$-th CBF constraints are active is given by
$\mathcal{S}_{i}=\{x\in\mathbb{R}^{n}\,|\,$ $f_{cl}(x)\!=\!f\!+\!\lambda_{i}G\nabla h_{i}-\lambda_{0}G\nabla V\}$ (33)
where $\lambda_{0},\lambda_{i}$ are the solutions of (32). The determinant of $C(x)$ is given by $|C(x)|=(p^{-1}+\lVert\nabla V\rVert_{G}^{2})\lVert\nabla h_{i}\rVert_{G}^{2}-\langle\nabla V,\nabla h_{i}\rangle_{G}^{2}$, which can be simplified by using the definition of $z_{2}$ in (27), yielding $|C(x)|=(p\eta)^{-1}\lVert\nabla h_{i}\rVert_{G}^{2}$. Notice that $|C(x)|\geq 0$, being zero if and only if $L_{g}h_{i}(x)=0$. Since we are considering the case $L_{g}h_{i}(x)\neq 0$, an expression for the inverse of $C(x)$ is then given by
$C(x)^{-1}$ $=\frac{p\eta}{\lVert\nabla h_{i}\rVert_{G}^{2}}\begin{bmatrix}\lVert\nabla h_{i}\rVert_{G}^{2}&\langle\nabla V,\nabla h_{i}\rangle_{G}\\ \langle\nabla V,\nabla h_{i}\rangle_{G}&p^{-1}+\lVert\nabla V\rVert_{G}^{2}\end{bmatrix}$ (34)
The derivative of (32) with respect to the $k$-th state component $x_{k}$ yields
$\partial_{k}C(x)\bar{\lambda}_{i}(x)\!+\!C(x)\partial_{k}\bar{\lambda}_{i}(x)=\partial_{k}b(x)$ (35)
where the operator $\partial_{k}$ denotes the partial derivative with respect to $x_{k}$. In the case where $L_{g}h(x)\neq 0$, $|C(x)|\neq 0$ and therefore the inverse (34) can be used to directly solve (35) for the partial derivatives of $\partial_{k}\bar{\lambda}_{i}(x)$:
$\partial_{k}\bar{\lambda}_{i}(x)$ $=C(x)^{-1}\underbrace{\left(\partial_{k}b(x)-\partial_{k}C(x)\bar{\lambda}_{i}(x)\right)}_{c(x)}$ (36)
To find expressions for $\partial_{k}\lambda_{0}(x)$, $\partial_{k}\lambda_{i}(x)$ using (36), expressions for $\partial_{k}C(x)$ and $\partial_{k}b(x)$ must be derived. Matrix $\partial_{k}C(x)$ is dependent on $\partial_{k}\lVert\nabla V\rVert_{G}$, $\partial_{k}\lVert\nabla h_{i}\rVert_{G}$ and $\partial_{k}\langle\nabla V,\nabla h_{i}\rangle_{G}$. Vector $\partial_{k}b(x)$ is dependent on $\partial_{k}F_{V}$ and $\partial_{k}F_{h_{i}}$. These can be computed using the derivatives
$\partial_{k}\langle u,v\rangle_{G}$ $\!=\!\langle\partial_{k}u,v\rangle_{G}\!+\!\langle\partial_{k}v,u\rangle_{G}\!+\!u^{\mathsf{T}}(\partial_{k}G)v$ (37)
$\partial_{k}F_{w}$ $\!=\!\langle\partial_{k}\nabla w,f\rangle\!+\!\langle\nabla w,\partial_{k}f\rangle\!+\!\beta^{\prime}(w)[\nabla w]_{k}$ (38)
with $u(x),v(x)\in\mathbb{R}^{n}$ replaced by $\nabla V(x)$ or $\nabla h_{i}(x)$ in (37), and with $w(x)\in\mathbb{R}$ replaced by $V(x)$ or $h_{i}(x)$ and $\beta$ replaced by $\gamma$ of class $\mathcal{K}$ or $\alpha$ of class $\mathcal{K}_{\infty}$ in (38). Using (37)-(38), an expression for $c(x)$ can be found:
$c(x)$ $=\begin{bmatrix}\langle\partial_{k}\nabla V,f_{cl}\rangle\!+\!\langle\nabla V,j_{k}\rangle\!+\!\gamma^{\prime}\!(V)\partial_{k}V\\ -\langle\partial_{k}\nabla h_{i},f_{cl}\rangle\!-\!\langle\nabla h_{i},j_{k}\rangle\!-\!\alpha^{\prime}\!(h_{i})\partial_{k}h_{i}\\ \end{bmatrix}$ (39)
$j_{k}$ $=\partial_{k}f+\lambda_{i}\partial_{k}(G\nabla h_{i})-\lambda_{0}\partial_{k}(G\nabla V)$
At a boundary equilibrium point $x_{e}\in\mathcal{E}_{\partial\mathcal{C}_{i}}$, $f_{cl}(x_{e})=0$ and $\lambda_{0}=p\gamma(V(x_{e}))$, simplifying (39) and allowing (36) to be written as
$\partial_{k}\bar{\lambda}_{i}(x_{e})$ $\!=\!C(x)^{-1}\!\begin{bmatrix}\nabla V^{\mathsf{T}}[J_{i}(x_{e},\lambda_{i})]_{k}\!+\!\gamma^{\prime}\!(V)\partial_{k}V\\ -\nabla h_{i}^{\mathsf{T}}[J_{i}(x_{e},\lambda_{i})]_{k}\!-\!\alpha^{\prime}\!(h_{i})\partial_{k}h_{i}\end{bmatrix}$ (40)
where $[J_{i}(x_{e},\lambda_{i})]_{k}$ denotes the $k$-th column of matrix $J_{i}(x_{e},\lambda_{i})$ defined at (30). Using $C(x_{e})^{-1}$ from (34), expressions for $\partial_{k}\lambda_{0}(x_{e})$ and $\partial_{k}\lambda_{i}(x_{e})$ follow from (40).
Equation (20) with $\lambda_{j}=0\,,\forall j\neq i$ gives the closed-loop system expression for Case 3, which is $f_{cl}(x)=f(x)-\lambda_{0}G(x)\nabla V(x)+\lambda_{i}G(x)\nabla h_{i}(x)$. Differentiating yields $\partial_{k}f_{cl}(x)=j_{k}-(\partial_{k}\lambda_{0})G\nabla V+(\partial_{k}\lambda_{i})G\nabla h_{i}$. At the boundary equilibrium point $x_{e}\in\mathcal{E}_{\partial\mathcal{C}_{i}}$, $\lambda_{0}=p\gamma(V(x_{e}))$ and $\partial_{k}f_{cl}(x_{e})$ can be written as
$\partial_{k}f_{cl}(x_{e})\!=\![J_{i}(x_{e},\lambda_{i})]_{k}$ $\!-\!\partial_{k}\lambda_{0}(x_{e})G\nabla V$
$\!+\!\partial_{k}\lambda_{i}(x_{e})G\nabla h_{i}$ (41)
Substituting the expressions for $\partial_{k}\lambda_{0}(x_{e})$ and $\partial_{k}\lambda_{i}(x_{e})$ obtained from (40) into (41) yields an involved expression that can be greatly simplified by using the definitions of $z_{1}$, $z_{2}$ in Definition IV.1, $\eta$ in (28) and property (iv) of $\eta$. After simplifications, the resulting expression for the $k$-th column of the closed-loop Jacobian at the boundary equilibrium point $x_{e}\in\mathcal{E}_{\partial\mathcal{C}_{i}}$ is
$\partial_{k}f_{cl}(x_{e})$ $\!=\!\left(\underbrace{I_{n}\!-\!Gz_{1}z_{1}^{\mathsf{T}}\!-\!p\eta Gz_{2}z_{2}^{\mathsf{T}}}_{I_{n}-GZN_{1}Z^{\mathsf{T}}}\right)\![J_{i}(x_{e},\lambda_{i})]_{k}$ (42)
$\qquad\qquad\qquad\qquad\qquad\quad-G[ZN_{1}\Psi Z^{\mathsf{T}}]_{k}$
Then, letting $k\in\{1,\cdots,n\}$ and combining the $n$ partial derivatives as column vectors in (42) to form the closed-loop Jacobian matrix $J_{cl}(x_{e})$ yields (29). $\square$
Lemma 2.
Assume $L_{g}h_{i}(x)\neq 0$, and consider orthogonality with respect to the inner product $\langle\cdot,\cdot\rangle_{G(x)}$.
  1. i
    If $L_{g}V(x)\neq 0$, define the set $\mathcal{Z}$ as $\mathcal{Z}=\{z_{1},z_{2}\}$.
  2. ii
    If $L_{g}V(x)=0$, define the set $\mathcal{Z}$ as $\mathcal{Z}=\{z_{1}\}$.
Let the set $\mathcal{W}=\{w_{1},\cdots,w_{\dim{\mathcal{W}}}\}$ be an orthonormal basis for $\text{span}\{\mathcal{Z}\}^{\perp}\subset\mathbb{R}^{n}$, that is, $\langle w_{i},w_{j}\rangle_{G}=\delta_{ij}$ for $i,j\in\{1,\cdots,\dim{\mathcal{W}}\}$ (since $\mathcal{W}$ is an orthonormal set) and $\langle w_{i},z_{k}\rangle_{G}=0$ for $i\in\{1,\cdots,\dim{\mathcal{W}}\}$, $k\in\{1,2\}$. Then, in both cases, the set $\mathcal{B}=\mathcal{Z}\cup\mathcal{W}$ is an orthogonal basis for $\mathbb{R}^{n}$.
Proof.
First notice that $\mathcal{B}=\mathcal{Z}\cup\mathcal{W}$ is an orthogonal set in both cases, since $\langle z_{1},z_{2}\rangle_{G(x)}=0$ and $\langle w_{i},z_{k}\rangle_{G(x)}=0$, $i\in\{1,\cdots,\dim{\mathcal{W}}\}$, $k\in\{1,2\}$. To prove that it is also a basis for $\mathbb{R}^{n}$, let the following be the linear independence equations for the vectors in $\mathcal{B}$ for $L_{g}V(x)\neq 0$ and $L_{g}V(x)=0$, respectively:
$\beta_{1}z_{1}+\beta_{2}z_{2}+\sum^{n-2}_{i=1}\beta_{i+2}w_{i}=0$ (43)
$\beta_{1}z_{1}+\sum^{n-1}_{i=1}\beta_{i+1}w_{i}=0$ (44)
Taking the inner product of (43)-(44) with $z_{1}$ yields $\beta_{1}=0$ in both cases, since $\langle w_{i},z_{1}\rangle_{G}=0\,,\forall i\in\{1,\cdots,\dim{\mathcal{W}}\}$. In case (i) where $L_{g}V(x)\neq 0$, $\dim{\mathcal{W}}=n-2$. Taking the inner product of (43) with $z_{2}$ yields $\beta_{2}=0$, since $z_{2}\neq 0$, $\langle z_{1},z_{2}\rangle_{G(x)}=0$ and $\langle w_{i},z_{2}\rangle_{G(x)}=0$, $i\in\{1,\cdots,\dim{\mathcal{W}}\}$. In case (ii) where $L_{g}V(x)=0$, $\dim{\mathcal{W}}=n-1$ and $z_{2}$ is identically zero, but it is also not contained in $\mathcal{Z}$. Notice that the absence of $z_{2}$ in $\mathcal{Z}$ is compensated by the presence of an extra basis vector for $\mathcal{W}$ appearing in the summation (since $\dim{\mathcal{W}}=n-1$ in this case). Taking the inner product of (43) or (44) with $w_{j}$, $j\in\{1,\cdots,\dim{\mathcal{W}}\}$ yields $\beta_{3}=\cdots=\beta_{\dim{\mathcal{W}}}=0$ or $\beta_{2}=\cdots=\beta_{\dim{\mathcal{W}}}=0$, respectively, for all terms in the summations, since $\langle z_{1},w_{j}\rangle_{G(x)}=\langle z_{2},w_{j}\rangle_{G(x)}=0$ and $\langle w_{i},w_{j}\rangle_{G(x)}=\delta_{ij}$. Therefore, the set $\mathcal{B}=\mathcal{Z}\cup\mathcal{W}$ forms a basis for $\mathbb{R}^{n}$ in the two considered cases. $\square$
Theorem 3 (Stability of Boundary Equilibria).
Under Assumption II.2, consider a boundary equilibrium point $x_{e}\in\mathcal{E}_{\partial\mathcal{C}_{i}}$ of the closed-loop system (3) with controller (4) such that $L_{g}h_{i}(x_{e})\neq 0$, with corresponding $i$-th KKT multiplier given by $\lambda_{e}\geq 0$. If there exists $v\in\{\nabla h_{i}(x_{e})\}^{\perp}$ such that
$v^{\mathsf{T}}J_{f_{i}}(x_{e},\lambda_{e})v>0\,,$ (45)
then $x_{e}$ is unstable. Otherwise, $x_{e}$ is stable. In particular, if $v^{\mathsf{T}}J_{f_{i}}(x_{e},\lambda_{e})v<0\,\,\forall v\in\{\nabla h_{i}(x_{e})\}^{\perp}$, then $x_{e}$ is asymptotically stable. The matrix function $J_{fi}$ is the Jacobian of the vector field $f_{i}$, as defined in (11).
Proof.
Consider a boundary equilibrium point $x_{e}\in\mathcal{E}_{\partial\mathcal{C}_{i}}$ with $L_{g}h_{i}(x_{e})\neq 0$. The corresponding Lyapunov equation for $x_{e}$ is then given by
$Y$ $=J_{cl}(x_{e})^{\mathsf{T}}X+XJ_{cl}(x_{e})$ (46)
where $J_{cl}(x_{e})$ is expressed by (29). Define
$X$ $=Z\Lambda_{z}Z^{\mathsf{T}}+W\Lambda_{w}W^{\mathsf{T}}>0$ (47)
where the columns of matrices $Z$ and $W$ are given by the basis vectors of $\mathcal{Z}$ and $\mathcal{W}$ of Lemma 2. Matrices $\Lambda_{z}$, $\Lambda_{w}$ are diagonal and positive definite, with dimensions combatible to the dimensions of $\mathcal{Z}$ and $\mathcal{W}$ from Lemma 2, that is, (i) if $L_{g}V(x_{e})\neq 0$, $\dim{\mathcal{Z}}=2$, $\dim{\mathcal{W}}=n-2$, (ii) if $L_{g}V(x_{e})=0$, $\dim{\mathcal{Z}}=1$, $\dim{\mathcal{W}}=n-1$. From the properties of vectors (26)-(27) and of the subspace $\mathcal{W}$, we have (i) $Z^{\mathsf{T}}GZ=diag\{1,p^{-1}(\eta^{-1}-1)\}$, (ii) $Z^{\mathsf{T}}GW=0$ and (iii) $W^{\mathsf{T}}GW=I_{n-2}$ if $L_{g}V(x_{e})\neq 0$ and $W^{\mathsf{T}}GW=I_{n-1}$ if $L_{g}V(x_{e})=0$. Substituting the closed-system Jacobian (29) and (47) in (46) and using the properties (i), (ii) and (iii) for matrices $Z$ and $W$, it is possible to write $Y$ in the following way, no matter the case considered ($L_{g}V(x_{e})\neq 0$ or $L_{g}V(x_{e})=0$):
$Y$ $=J_{i}^{\mathsf{T}}\overline{X}+\overline{X}J_{i}-Z\Omega Z^{\mathsf{T}}$ (48)
where $\Omega=\Omega^{\mathsf{T}}\geq 0$. The expressions for $\overline{X}$, $\Omega$ and $Z$ depend on the considered case:
Case (i): if $L_{g}V(x_{e})\neq 0$, $\dim{\mathcal{Z}}=2$, $Z=[\,z_{1}\,\,\,z_{2}\,]\in\mathbb{R}^{n\times 2}$, $\Lambda_{z}=\text{diag}\{\lambda_{z_{1}},\lambda_{z_{2}}\}$. In (48), $\overline{X}=\eta\lambda_{z_{2}}z_{2}z_{2}^{\mathsf{T}}+W\Lambda_{w}W^{\mathsf{T}}\geq 0$ and $\Omega=\Psi^{\mathsf{T}}N\Lambda_{z}+\Lambda_{z}N\Psi$, with $N=\text{diag}\{1,1-\eta\}$. Here, $W=[\,w_{1}\,\,\cdots\,\,w_{n-2}\,]\in\mathbb{R}^{n\times(n-2)}$.
Case (ii): if $L_{g}V(x_{e})=0$, $\dim{\mathcal{Z}}=1$, $Z=z_{1}\in\mathbb{R}^{n}$, $\Lambda_{z}=\lambda_{z_{1}}\in\mathbb{R}_{+}$. In (48), $\overline{X}=W\Lambda_{w}W^{\mathsf{T}}\geq 0$ and $\Omega=2\lambda_{z_{1}}\alpha^{\prime}(h_{i})$. Here, $W=[\,w_{1}\,\,\cdots\,\,w_{n-1}\,]\in\mathbb{R}^{n\times(n-1)}$.
In both cases, $\overline{X}$ has no term $z_{1}z_{1}^{\mathsf{T}}$. By Chetaev’s instability theorem, $x_{e}$ is unstable if there exists $v\in\mathbb{R}^{n}$ such that $v^{\mathsf{T}}Yv>0$ in (48) [4]. Then, the quadratic form $v^{\mathsf{T}}Yv$ yields
$v^{\mathsf{T}}Yv=2v^{\mathsf{T}}\overline{X}J_{i}v-v^{\mathsf{T}}Z\Omega Z^{\mathsf{T}}v$ (49)
Let $v\in\{\nabla h_{i}(x_{e})\}^{\perp}$. Then:
Case (i): if $L_{g}V(x_{e})\neq 0$, the second term on the right-hand side of (49) becomes
$v^{\mathsf{T}}Z\Omega Z^{\mathsf{T}}v$ $=v^{\mathsf{T}}\underbrace{\left(\gamma^{\prime}(V)\sigma_{z_{2}}(1-\eta)z_{2}z_{2}^{\mathsf{T}}\right)}_{p\gamma^{\prime}(V)\overline{X}Gz_{2}z_{2}^{\mathsf{T}}}v$ (50)
$=v^{\mathsf{T}}\left(p\gamma^{\prime}(V)\overline{X}G\nabla V\nabla V^{\mathsf{T}}\right)v$ (51)
where we have used the fact that $\langle w_{i},z_{2}\rangle_{G}=0$ and property (iii) of (28). Then, (49) can be rewritten as
$v^{\mathsf{T}}Yv=2v^{\mathsf{T}}\overline{X}\underbrace{\left(J_{i}-p\gamma^{\prime}(V)G\nabla V\nabla V^{\mathsf{T}}\right)}_{J_{f_{i}}(x_{e},\lambda)}v$ (52)
Case (ii): if $L_{g}V(x_{e})=0$, the second term on the right-hand side of (49) vanishes, since $Z=z_{1}$ and therefore, $Z^{\mathsf{T}}v=0$. Then, (49) yields
$v^{\mathsf{T}}Yv=2v^{\mathsf{T}}\overline{X}J_{i}v=2v^{\mathsf{T}}\overline{X}\left(J_{f_{i}}(x_{e},\lambda)\big{|}_{G\nabla V=0}\right)v$ (53)
Therefore, in both cases (52)-(53) result in $v^{\mathsf{T}}Yv=2v^{\mathsf{T}}\overline{X}J_{f_{i}}(x_{e},\lambda)v$. Given any matrix $M$, since $\overline{X}$ is symmetric and positive semi-definite, matrices $\overline{X}M$ and $\overline{X}^{\frac{1}{2}}M\overline{X}^{\frac{1}{2}}$ share the same nonzero eigenvalues. Since the spectra of $\overline{X}^{\frac{1}{2}}M\overline{X}^{\frac{1}{2}}$ is real, so is the spectra of $\overline{X}M$. Since $\overline{X}\geq 0$ is arbitrary, with its one-dimensional nullspace spanned by $G\,z_{1}\neq 0$, it is always possible to choose $\Lambda_{z}$ and $\Lambda_{w}$ in such a way that $v\in\{\nabla h_{i}(x_{e})\}^{\perp}$ is an eigenvector of $\overline{X}$ with a corresponding strictly positive eigenvalue $\lambda(\overline{X})>0$. Then, $v^{\mathsf{T}}Yv$ yields
$v^{\mathsf{T}}Yv=2\lambda(\overline{X})v^{\mathsf{T}}J_{f_{i}}(x_{e},\lambda)v$ (54)
Then, $x_{e}\in\mathcal{E}_{\partial\mathcal{C}_{i}}$ is unstable if the right-hand side of (54) is strictly positive, demonstrating (45).
To show that $x_{e}\in\mathcal{E}_{\partial\mathcal{C}_{i}}$ is locally stable otherwise, we proceed as follows. The first order Taylor series approximation of the closed-loop system on a neighborhood of $x_{e}$ is $\dot{x}=J_{cl}(x_{e})\Delta x$ with $\Delta x=(x-x_{e})$ being a disturbance vector around the equilibrium point. Let us write this disturbance vector using the basis $\{z_{1}(x_{e}),v_{1},\cdots,v_{n-1}(x_{e})\}$, where the $v_{1},\cdots,v_{n-1}$ are fixed basis vectors for $\{\nabla h_{i}(x_{e})\}^{\perp}$. Therefore, $\Delta x=\beta z_{1}(x_{e})+v$, with $v=\sum^{n-1}_{i=1}\beta_{i}v_{i}$. Note that $v^{\mathsf{T}}\nabla h_{i}(x_{e})=0$ by construction. Here, $\beta,\beta_{1},\cdots,\beta_{n-1}$ represent the coordinates of $\Delta x$ in the new basis. Computing the inner product $\langle z_{1}(x_{e}),\dot{x}\rangle$ yields
$\langle z_{1}(x_{e}),\dot{x}\rangle$ $=\langle z_{1},\dot{\beta}z_{1}+\sum^{n}_{i=1}\dot{\beta_{i}}v_{i}\rangle=\dot{\beta}\lVert z_{1}(x_{e})\rVert^{2}$
$=z_{1}^{\mathsf{T}}J_{cl}(x_{e})(\beta z_{1}(x_{e})+v)$ (55)
Since $z_{1}^{\mathsf{T}}J_{cl}(x_{e})=-\alpha^{\prime}(h_{i}(x_{e}))z_{1}(x_{e})^{\mathsf{T}}$ in (55), the dynamics of $\beta$ is given by $\dot{\beta}=-\alpha^{\prime}(h_{i}(x_{e}))\beta$. Since $\alpha^{\prime}$ is a $\mathcal{K}_{\infty}$ function, $\alpha^{\prime}(h_{i}(x_{e}))>0$, which means that $\beta\rightarrow 0$. Replacing the dynamics of $\beta$ into the Taylor expansion yields the following dynamics for $v$:
$\dot{v}$ $=\beta\left(J_{cl}(x_{e})+\alpha^{\prime}(h_{i})I_{n}\right)z_{1}+J_{cl}(x_{e})v$ (56)
Define the Lyapunov candidate $V(\beta,v)=\frac{1}{2}\beta^{2}+v^{\mathsf{T}}Xv>0$, with $X$ given by (47). Taking its time derivative and using the dynamics of $\beta$ and (56) yields
$\dot{V}$ $=-\alpha^{\prime}(h_{i})\beta^{2}+v^{\mathsf{T}}Yv$ (57)
$+2v^{\mathsf{T}}X\left(J_{cl}(x_{e})+\alpha^{\prime}(h_{i})I_{n}\right)z_{1}(x_{e})\beta$
Equation (57) shows that, since the dynamics of $\beta$ is decoupled and asymptotically stable, the sign of $\dot{V}$ is eventually determined by the term $v^{\mathsf{T}}Yv$. By (54), if $v^{\mathsf{T}}J_{f_{i}}(x_{e},\lambda)v\leq 0\,\,\forall v\in\{\nabla h_{i}(x_{e})\}^{\perp}$, then $v^{\mathsf{T}}Yv\leq 0$, and $x_{e}$ is stable. In particular, if $v^{\mathsf{T}}J_{f_{i}}(x_{e},\lambda)v<0\,\,\forall v\in\{\nabla h_{i}(x_{e})\}^{\perp}$, then $x_{e}$ is asymptotically stable. $\square$
V CLF Compatibility
In light of Theorem 3, the stability properties of boundary equilibrium points with $L_{g}h(x)\neq 0$ are determined by the Jacobian $J_{f_{i}}$ in (11), which depends on the system dynamics and on the geometry of the CLF and CBFs. Since the system dynamics must be kept general and the CBFs must provide a model for the safety requirements of the problem, we consider the following question: is it possible to find a valid CLF such that all boundary equilibrium points that are unremovable by the technique of [9] are either removed or unstable? Our objective in this section is to prove that this is indeed possible by providing a method for computing such CLF, considering a particular type of system and CLF-CBF geometry.
Definition V.1 (CLF Compatibility).
Under Assumptions II.1-III.1, a CLF with global minimum in $x_{0}\in\mathbb{R}^{n}$ is said to be $i$-th compatible if $x_{0}$ is the only stable equilibrium point of the closed-loop system (3) with only the $i$-th CBF implemented in the QP controller (4).
Under Assumptions II.1-III.1, it follows immediately that the CLF minimum $x_{0}$ is quasi-globally asymptotically stable for the closed-loop system (3) with the QP controller (4) implemented with only the $i$-th CBF constraint and an $i$-th compatible CLF $V$. Even when boundary equilibrium points exist, since they are all unstable, the trajectories fail to converge to the origin only in a set of measure zero (that is, on the existing boundary equilibrium points).
V-A Quadratic CLF Compatibility
Consider the following classes of systems:
  1. (i)
    Nonlinear systems of the form $\dot{x}=g(x)u$ with full-rank $g(x)$ for all $x\in\mathbb{R}^{n}$.
  2. (ii)
    Linear Time-Invariant systems $\dot{z}=Az+Bu$, $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$ (LTI), where $z=x-x_{0}$.
Let the transformed CLF $\overline{V}$ and the $N$ CBFs be parametrized by quadratic polynomials on $\mathbb{R}^{n}$, as
$\overline{V}(x)$ $=0.5\Delta x^{\mathsf{T}}H_{\overline{V}}\Delta x\geq 0\,,$ $\Delta x=x-x_{0}$ (58)
$h_{i}(x)$ $=0.5\left(\Delta x_{i}^{\mathsf{T}}H_{h_{i}}\Delta x_{i}-1\right)\,,$ $\Delta x_{i}=x-c_{i}$ (59)
where the parameters of the CLF and $i$-th CBF are: the constant Hessian matrices $H_{\overline{V}},H_{h_{i}}\in\mathcal{S}_{+}$ determining the elliptical shapes of their level sets and their centers $x_{0},x_{i}\in\mathbb{R}^{n}$, $i\in\{0,1,\cdots,N\}$. Recall that due to property (ii) of Proposition 1, $V$ can be computed from $\overline{V}$ as defined in (58), and then used in the QP controller (4).
Using the gradients of (58)-(59), consider the expressions for $f_{i}(x,\lambda)$ for each class of systems after performing a state translation $\nu=\Delta x_{i}$ ($x=\nu+c_{i}$):
Case (i) For nonlinear systems of the form $\dot{x}=g(x)u$ with full-rank $g(x)$, $f_{i}(\nu,\lambda)=G(\nu)(P(\lambda)\nu-w)$, where $G(x)>0$ and $P(\lambda)=\lambda M-N$ is a symmetric Linear Matrix Pencil (LMP) [12] with $M=H_{h_{i}}$, $N=pH_{\overline{V}}$.
Case (ii) For LTI systems, $f_{i}(\nu,\lambda)=P(\lambda)\nu-w$, where $P(\lambda)=\lambda M-N$ is a regular LMP with $M=BB^{\mathsf{T}}H_{h_{i}}$, $N=pBB^{\mathsf{T}}H_{\overline{V}}-A$. In both cases, $w=N(x_{i}-x_{0})$.
Define the scalar function $q(\nu)=\nu^{\mathsf{T}}H_{h_{i}}\nu$ associated to the $i$-th CBF. Then, in both cases, the boundary equilibrium points on $\partial\mathcal{C}_{i}$ must satisfy
$P(\lambda)\nu$ $=w\,,\quad\lambda>0$ (60)
$q(\nu)$ $=1$ (61)
The solutions of (60)-(61) are connected to the theory of LMPs [12]. Let $\sigma_{P}=\{\lambda\in\mathbb{C}:\det P(\lambda)=0\}$ be the spectrum of the pencil $P$. Since $P(\lambda)$ is regular, its inverse matrix $P(\lambda)^{-1}$ exists $\forall\lambda\not\in\sigma_{P}$. Then, equation (60) can be solved for $\nu$, yielding
$\nu(\lambda)=P(\lambda)^{-1}w$ (62)
Equation (62) describes the equilibrium manifold $f_{i}(\nu,\lambda)=0$ where boundary equilibrium points occur. Substituting it into $q(\nu)=\nu^{\mathsf{T}}H_{h_{i}}\nu$ yields
$q(\lambda)$ $=w^{\mathsf{T}}P(\lambda)^{-\mathsf{T}}H_{h_{i}}P(\lambda)^{-1}w=\frac{n(\lambda)}{|P(\lambda)|^{2}}$ (63)
where $n(\lambda)=w^{\mathsf{T}}\operatorname{adj}{P(\lambda)}^{\mathsf{T}}H_{h_{i}}\operatorname{adj}{P(\lambda)}w$ and $|P(\lambda)|^{2}$ are non-negative polynomials with coefficients in $\mathbb{R}$. Equation (63) is defined as the Q-function for the $i$-th barrier $h_{i}$. For the classes of systems described in Section V-A, it encodes all the necessary information for computing the equilibrium points on the $i$-th boundary. Note that, due to (61), every $\lambda_{e}\geq 0\notin\sigma_{P}$ satisfying $q(\lambda_{e})=1$ corresponds to a non-degenerate equilibrium solution, where the corresponding boundary equilibrium point is $x_{e}=\nu(\lambda_{e})+c_{i}=P(\lambda_{e})^{-1}w+c_{i}\in\partial\mathcal{C}_{i}$.
Theorem 4 (Q-Function Properties).
Consider the safety-critical control problem described in Section V-A, under Assumptions II.1-III.1, and the Q-function $q(\lambda)$ associated to the $i$-th CBF.
(i) If Assump. II.1 holds, then $q(0)\geq 1$.
(ii) If $q(\lambda)$ is proper, the closed-loop system (3) has at least one boundary equilibrium point.
(iii) Let $R(\lambda)\in\mathbb{R}^{n\times(n-1)}$ be a matrix polynomial on the orthogonal space of $H_{h_{i}}\operatorname{adj}{P(\lambda)}w\in\mathbb{R}^{n}$, that is, satisfying $R(\lambda)^{\mathsf{T}}H_{h_{i}}\operatorname{adj}{P(\lambda)}w=0\,\,\forall\lambda\in\mathbb{R}$. Define the stability matrix polynomial as
$S(\lambda)=R(\lambda)^{\mathsf{T}}\left(P(\lambda)+P(\lambda)^{\mathsf{T}}\right)R(\lambda)\,.$ (64)
Then, an equilibrium point $x_{e}=\nu(\lambda_{e})+x_{i}\in\partial\mathcal{C}_{i}$ is stable if and only if $S(\lambda_{e})\leq 0$. Otherwise, it is unstable.
(iv) The maximum number of negative semi-definite intervals of $S(\lambda)$ is $n$.
Proof.
For $P(\lambda)$ defined in both Cases (i) and (ii), evaluating (60) at $\lambda=0$ yields $P(0)\nu=w$, yielding $\nu(0)=x_{0}-c_{i}$. Therefore, $q(0)=(x_{0}-c_{i})^{\mathsf{T}}H_{h_{i}}(x_{0}-c_{i})$. Then, by (59), $q(0)>1$ is equivalent to $h_{i}(x_{0})>0$, which means that $x_{0}\in\mathcal{C}_{i}$, that is, Assump. II.1 is satisfied. This proves (i).
Consider an arbitrary closed interval $\mathcal{I}\subset\mathbb{R}_{+}$. If $q(\lambda)>1$ for all $\lambda\in\mathcal{I}$, then $\mathcal{I}$ does not contain equilibrium point solutions. Using (63), $q(\lambda)>1\Rightarrow n(\lambda)-|P(\lambda)|^{2}>0$ over $\mathcal{I}$. If it was possible to guarantee this condition for the entire positive real line $\mathcal{I}=\mathbb{R}_{+}$, then no boundary equilibrium points would exist. However, this is impossible in general, since $q(\lambda)\geq 0$ and $\lim_{\lambda\rightarrow+\infty}q(\lambda)=0$ in case $q(\lambda)$ is proper (which happens if the pencil $P$ has no generalized eigenvalues at infinity). For the considered problem, this demonstrates the impossibility of removing all undesirable equilibrium points for certain types of systems. This proves (ii).
Next, consider the translated state $\nu=x-c_{i}$ and the CL and LTI systems in Section V-A. Then, the boundary equilibrium point is $\nu_{e}=\nu(\lambda_{e})\in\partial\mathcal{C}_{i}$ for some $\lambda_{e}\geq 0$ such that $q(\lambda_{e})=1$, and $\nu$ as defined in (62).
Case (i) For CL systems, the columns of $J_{f_{i}}(\nu,\lambda)$ are given by $\partial_{k}f_{i}(\nu,\lambda_{e})=\partial_{k}G(\nu)(P(\lambda)\nu-w)+G(\nu)[P(\lambda)]_{k}$. Using (60), $J_{f_{i}}(\nu_{e},\lambda_{e})=G(\nu_{e})P(\lambda_{e})$.
Case (ii) For LTI systems, the columns of $J_{f_{i}}(\nu,\lambda)$ are given by $\partial_{k}f_{i}(\nu,\lambda)=[P(\lambda)]_{k}$. Using (60), $J_{f_{i}}(\nu_{e},\lambda)=P(\lambda_{e})$.
Since $\nabla h_{i}(\nu)=H_{h_{i}}\nu$, using (62), the vector rational function $\nabla h_{i}(\lambda)=H_{h_{i}}P(\lambda)^{-1}w$ describes the barrier gradient in the equilibrium manifold consisting of states $\nu$ such that $f_{i}(\nu,\lambda)=0$. At the equilibrium point $\nu_{e}$, we have $\nabla h_{i}(\nu_{e})=\nabla h_{i}(\lambda_{e})$. From Theorem 3, for any of the two cases, if there exists $v\in\{\nabla h_{i}(\lambda_{e})\}^{\perp}$ such that $v^{\mathsf{T}}P(\lambda_{e})v>0$, then $\nu_{e}$ is unstable. Let $v=\mathsf{P}_{\nabla h_{i}}(\lambda_{e})z$ with $z\in\mathbb{R}^{n}$, where $\mathsf{P}_{\nabla h_{i}}(\lambda)=\lVert\nabla h_{i}(\lambda)\rVert^{2}I_{n}-\nabla h_{i}(\lambda)\nabla h_{i}(\lambda)^{\mathsf{T}}$. Then, $v$ is a projection into $\{\nabla h_{i}(\nu_{e})\}^{\perp}$. Substituting $v$ into (45), $\nu_{e}$ is unstable if $\exists z\in\mathbb{R}^{n}$ such that
$z^{\mathsf{T}}S_{null}(\lambda_{e})z>0\,,$ (65)
where $S_{null}(\lambda)=\mathsf{P}_{\nabla h_{i}}(\lambda)^{\mathsf{T}}(P(\lambda)+P(\lambda)^{\mathsf{T}})\mathsf{P}_{\nabla h_{i}}(\lambda)$. By construction, $\nabla h_{i}(\lambda)$ is in the null-space of $S_{null}(\lambda)$ for all $\lambda\in\mathbb{R}$.
The null-space of $\nabla h_{i}(\lambda)$ is the same as that of the vector polynomial $H_{h_{i}}\operatorname{adj}{P(\lambda)}w$, since they differ only by a scalar factor of $|P(\lambda)|^{-1}$. Since $|P(\lambda)|$ is of maximum degree $n$ and $P(\lambda)\operatorname{adj}{P(\lambda)}=|P(\lambda)|I_{n}$, $\operatorname{adj}{P(\lambda)}$ is a polynomial matrix of maximum degree $n-1$. Let $H_{h_{i}}\operatorname{adj}{P(\lambda)}w=v_{0}+\lambda v_{1}+\cdots+\lambda^{l}v_{l}$, $l\leq n-1$, with $v_{i}$ being constant vector coefficients. Let $r(\lambda)=r_{0}+\lambda r_{1}+\cdots+\lambda^{d}r_{d}\in\mathbb{R}^{n}$ be a vector polynomial of degree $d\in\mathbb{N}$ in the null-space of $\nabla h_{i}(\lambda)$. Since $r(\lambda)^{\mathsf{T}}(v_{0}+\lambda v_{1}+\cdots+\lambda^{l}v_{l})=0\,\,\forall\lambda\in\mathbb{R}$, their coefficients must satisfy $\sum^{k}_{i=0}v_{i}^{\mathsf{T}}r_{i-k}=0$ for $k=\{0,l+d\}$. These $l+d+1$ equations can be stacked in matrix form $V\overline{r}=0$, with $\overline{r}^{\mathsf{T}}=[\,r_{0}^{\mathsf{T}}\,\,r_{1}^{\mathsf{T}}\,\,\cdots\,\,r_{d}^{\mathsf{T}}\,]\in\mathbb{R}^{(d+1)n}$ , $V\in\mathbb{R}^{(l+d+1)\times(d+1)n}$ [13][Chapter XII, Section 3]. Since $\mathbb{R}^{n}=\text{span}\{H_{h_{i}}\operatorname{adj}{P(\lambda)}w\}\bigoplus\{H_{h_{i}}\operatorname{adj}{P(\lambda)}w\}^{\perp}\,,\forall\lambda\in\mathbb{R}$ (a direct sum of orthogonal subspaces), one can always obtain $n-1$ linearly independent basis vectors for $\{H_{h_{i}}\operatorname{adj}{P(\lambda)}w\}^{\perp}$ from the vector coefficients drawn from a basis of $\mathcal{N}(V)$.
Then, the arbitrary vector $z$ from (65) can be decomposed using the basis
$\{\nabla h_{i}(\lambda_{e}),r_{1}(\lambda_{e}),\cdots,r_{n-1}(\lambda_{e})\}$ (66)
where $r_{i}(\lambda)$, $i\in\{1,\cdots,n-1\}$ are basis polynomials for $\{H_{h_{i}}\operatorname{adj}{P(\lambda)}w\}^{\perp}$. Then, $z=\beta\nabla h_{i}(\lambda_{e})+R(\lambda_{e})\alpha$, where $R(\lambda)=[\,r_{1}(\lambda)\,\,\cdots\,\,r_{n-1}(\lambda)\,]\in\mathbb{R}^{n\times(n-1)}$ is a matrix polynomial of degree $d$ whose columns are in $\{H_{h_{i}}\operatorname{adj}{P(\lambda)}w\}^{\perp}$, and $\beta\in\mathbb{R}$, $\alpha\in\mathbb{R}^{n-1}$ are the coordinates of $z$ in the basis (66). Then, we have $\mathsf{P}_{\nabla h_{i}}(\lambda)z=\lVert\nabla h(\lambda)\rVert^{2}R(\lambda)\alpha$, and the left side of (65) becomes
$z^{\mathsf{T}}S_{null}(\lambda_{e})z\!=\!\lVert\nabla h(\lambda_{e})\rVert^{4}\alpha^{\mathsf{T}}S(\lambda_{e})\alpha$ (67)
with $S(\lambda)$ as defined in (64). Since $\alpha$ is arbitrary, from Theorem 3, we conclude that $\nu_{e}$ is stable if and only if $S(\lambda)$ is negative semi-definite at $\lambda_{e}$. Otherwise, $\exists\alpha\in\mathbb{R}^{n-1}$ such that (67) is strictly positive, and $\nu_{e}$ is unstable. This proves (iii).
The polynomial matrix $S(\lambda)\in\mathbb{R}^{n-1\times n-1}$ has maximum degree 3 (odd) and its leading coefficient is positive semi-definite. That means that there exist threshold values $\sigma_{+}$, $\sigma_{-}$ such that $S(\lambda)\geq 0$ for all $\lambda\geq\sigma_{+}$ and $S(\lambda)\leq 0$ for all $\lambda\leq\sigma_{-}$, respectively. Furthermore, its determinant $|S(\lambda)|$ has a maximum of $3(n-1)$ real roots, which are exactly the values of $\lambda$ where the eigenvalue curves of $S(\lambda)$ change sign. That means that all $n-1$ eigenvalue curves of $S(\lambda)$ must go from negative to positive as $\lambda$ increases. Then, excluding these $n-1$ roots, a total of $2(n-1)$ roots remain, which can result in a maximum of $n-1$ negative semi-definite intervals for $S(\lambda)$ (in case all roots of $|S(\lambda)|$ are real), plus the negative semi-definite interval of infinite length $\mathcal{I}_{1}=(-\infty,\sigma_{-}]$. Therefore, in the worst case, $n$ negative semi-definite intervals for $S(\lambda)$ exist. This proves (iv). $\square$
Corollary 1.
Under Assumptions II.1-III.1 and considering LTI or drift-less full-rank systems, a quadratic CLF (58) is $i$-th compatible if and only if $S(\lambda_{e})$ is not negative semi-definite at the positive real roots $\lambda_{e}\geq 0\in\mathbb{R}$ of the polynomial $z(\lambda)=n(\lambda)-|P(\lambda)|^{2}$.
Proof.
This result follows directly from property (iii) of Theorem 4. Notice that the positive roots of $z(\lambda)=n(\lambda)-|P(\lambda)|^{2}$ correspond to the equilibrium solutions $q(\lambda)=\frac{n(\lambda)}{|P(\lambda)|^{2}}=1$. If all of them occur at the regions where $S(\lambda)$ is not negative semi-definite, then these roots correspond to unstable boundary equilibrium points. By Assump. III.1, no interior equilibrium points other then the CLF minimum exist. Then, we conclude that the CLF (58) is $i$-th compatible. $\square$
Figure 1: Example of a Q-function in two dimensions.
Figure 1 shows the graphs of $q(\lambda)$ and $z(\lambda)=n(\lambda)-|P(\lambda)|^{2}$ for the LTI system $\dot{x}_{k}=-2x_{k}+u_{k}$, $k=1,2$, CLF $V(x)=2.5x_{1}^{2}+4.0x_{2}^{2}$, centered on $x_{0}=(0,0)$ and CBF $h_{1}(x)=1.95(x_{1}-6)^{2}-0.61(x_{1}-6)x_{2}+0.28x_{2}^{2}-0.5$, centered on $c_{1}=(6,0)$. The asymptotes of $q(\lambda)$ occur at the two generalized eigenvalues of the pencil $P$. Notice that $q(0)\geq 1$, or equivalently $z(0)\geq 0$. From Theorem 4(i), this implies Assump. II.1, that is, $x_{0}\in\mathcal{C}$, which is indeed true from the CBF expression, since $h(0)\geq 0$. In this example, $S(\lambda)$ is simply a $1\times 1$ matrix polynomial (a scalar polynomial) of degree $3$ with real coefficients. Therefore, $|S(\lambda)|$ has three roots in $\mathbb{C}$. Theorem 4(iv) implies that a maximum of $n=2$ negative semi-definite intervals can occur for $S(\lambda)$: in this example, $S(\lambda)\leq 0$ in the interval (,23)($\infty$,$\approx$ 23)( - ∞ , ≈ 23 ) (red strip) and $S(\lambda)\geq 0$ in the interval $(\approx 23,+\infty)$ (blue strip). These intervals are separated by the real root of $|S(\lambda)|$ at $\lambda\approx 23$ (star-shaped dot); the remaining two roots of $|S(\lambda)|$ are complex-conjugates. There is one solution $z(\lambda_{e})=0$ around $\lambda_{e}\approx 16$ corresponding to an stable equilibrium point (red dot) and other two around $\lambda_{e}\approx 28,42$ corresponding to unstable equilibrium points (blue dots). From Corollary 1, the CLF is compatible with $h_{1}$.
Theorem 5 (Compatibility Barrier).
Under Assumptions II.1-III.1, consider the Q-function $q(\lambda)$ associated to the $i$-th CBF. Let $\mathcal{I}=(-\infty,\sigma_{-}]$ be the first negative semi-definite interval of $S(\lambda)$, that is, $S(\lambda)\leq 0\,\,\forall\lambda\leq\sigma_{-}$. Then, the CLF (58) is $i$-th compatible if
$B(q)=\min_{\lambda\in\mathcal{I}\cap\mathbb{R}_{\geq 0}}n(\lambda)-\epsilon|P(\lambda)|^{2}\geq 0$ (68)
$S^{\prime}(\lambda)\geq 0$ (69)
where $\epsilon>1$.
Proof.
In condition (68), $B(q)$ represents a barrier function for the semi-definite interval $\mathcal{I}$: that is, under Assump. II.1-II.2, if $B(q)$ is non-negative, then there are no roots of $z_{\epsilon}(\lambda)$ in $\mathcal{I}\cap\mathbb{R}_{\geq 0}$. Then, $q(\lambda)=\frac{n(\lambda)}{|P(\lambda)|^{2}}\geq\epsilon>1$ in $\mathcal{I}\cap\mathbb{R}_{\geq 0}$, which means that no boundary equilibrium solutions of (60)-(61) exist in $\mathcal{I}$. Condition (69) ensures that the eigenvalues of $S(\lambda)$ are monotonically increasing. This is a sufficient condition to ensure that no negative semi-definite interval of $S(\lambda)$ other than $\mathcal{I}$ exists, and therefore equilibrium solutions of (60)-(61) occurring in $\mathbb{R}_{\geq 0}\setminus\mathcal{I}$ correspond to unstable equilibrium points. Under Assump. III.1, no interior equilibrium points other than $x_{0}$ exist. This shows that no stable equilibrium other than $x_{0}$ exists under conditions (68)-(69). Thus, the CLF is $i$-th compatible. $\square$
Remark V.1.
As a functional of the Q-function associated to the $i$-th CBF, the barrier function $B(q)$ in (68) is dependent on the system dynamics, CLF and $i$-th CBF geometry. Furthermore, Theorem 5 is a sufficient, although not necessary condition for CLF $i$-th compatibility.
V-B Compatible CLF Controller
In this section, our objective is twofold: (i) to propose a “compatibilization” algorithm for computing a compatible CLF from a non-compatible one, and (ii) to propose a control strategy to smoothly transform the CLF used in the QP-controller (4) towards the compatible CLF computed from the compatibilization algorithm, in the regions of the state space where boundary equilibrium points occur.
Definition V.2.
Hessian $H$ is $i$-th compatible if its corresponding CLF $V(x)=\frac{1}{2}\Delta x^{\mathsf{T}}H\Delta x$ is $i$-th compatible.
Let $\overline{V}_{r}(x)=\frac{1}{2}\Delta x^{\mathsf{T}}H_{\overline{V}_{r}}\Delta x$ be a reference quadratic CLF centered on $x_{0}\in\mathcal{C}$ (Assump. II.1 holds) and define the following optimization problem related to the $i$-th CBF:
$H_{\overline{V}_{i}}=$ $\operatorname*{argmin}_{H\in\mathcal{S}^{n}_{+}}\,\,\lVert H-H_{\overline{V}_{r}}\rVert_{\mathcal{F}}^{2}$ (70)
[…]
$\qquad\,\,\,B(q)\geq 0$ (compatibility)
$\qquad\,\,\,S^{\prime}(\lambda)\geq 0$ (monotonicity)
The result of optimization (70) is the Hessian $H_{\overline{V}_{i}}$ of the closest quadratic $\overline{V}_{i}(x)=\frac{1}{2}\Delta x^{\mathsf{T}}H_{\overline{V}_{i}}\Delta x$ to the reference CLF $\overline{V}_{r}$ satisfying:
(i) $\overline{V}_{i}$ is a valid CLF since it satisfies the CLF condition for the given LTI system: $L_{f}\overline{V}_{i}\leq 0\rightarrow H_{\overline{V}_{i}}A+A^{\mathsf{T}}H_{\overline{V}_{i}}\leq 0$. For driftless systems, the CLF condition is always satisfied since $f(x)=0$.
(ii) $\overline{V}_{i}$ is $i$-th compatible (due to Theorem 5). Likewise, $H_{\overline{V}_{i}}$ is $i$-th compatible.
Thus, if the reference $H_{\overline{V}_{r}}$ is already $i$-th compatible, the result of (70) is $H_{\overline{V}_{i}}=H_{\overline{V}_{r}}$.
Remark V.2.
In optimization (70), the barrier $B(q)$ depends on polynomials such as $n(\lambda)$, $|P(\lambda)|^{2}$ and $S(\lambda)$, which can be efficiently computed using computational methods for polynomial manipulation. These polynomials depend on the system and CLF-CBF parameters. Therefore, $B(q)$ is dependent on the optimization variable $H$, and must be recomputed at each solver iteration. Any solver supporting non-convex constrained optimization could be used, such as Sequential Least Squares Programming (SLSQP) [14].
Let $\{H_{\overline{V}_{1}},\cdots,H_{\overline{V}_{N}}\}$ be a set of $N$ compatible Hessians computed using (70), where $H_{\overline{V}_{i}}$ is the closest $i$-th compatible Hessian to the reference $H_{\overline{V}_{r}}$. Define a parametric CLF $\overline{V}(x,\pi)=\frac{1}{2}\Delta x^{\mathsf{T}}H_{\overline{V}}(\pi)\Delta x$ with parametrized Hessian given by
$H_{\overline{V}}(\pi)=L(\pi)^{\mathsf{T}}L(\pi)\in\mathcal{S}^{n}_{\geq 0}$ (71)
where $\pi\in\mathbb{R}^{\dim{\mathcal{S}^{n}}}$ is a state vector defining the geometry of the level sets of $\overline{V}$. We seek to design a controller for $\pi$, so that the level sets of $\overline{V}$ are dynamically changed. Using an integrator as the CLF shape state dynamics $\dot{\pi}=u_{\pi}$, define a Lyapunov function candidate as
$V_{\pi}(\pi,H_{r})=\frac{1}{2}\lVert H_{\overline{V}}(\pi)-H_{r}\rVert_{\mathcal{F}}^{2}$ (72)
where $H_{r}\in\mathcal{S}^{n}_{+}$ is a constant Hessian. If a stabilization controller for $\dot{\pi}=u_{\pi}$ is designed in such a way that $\dot{V}_{\pi}=\nabla V_{\pi}^{\mathsf{T}}u_{v}\leq 0$, then $H_{\overline{V}}(\pi)$ approaches $H_{r}$. This way, the level sets of $\overline{V}$ are smoothly adapted to match the level sets of a CLF with Hessian $H_{r}$ and center $x_{0}$.
Now, consider the QP controller (4) with the CLF $V$ computed from the inverse CLF transformation of $\overline{V}(x,\pi)$, as defined in Definition II.4. This transformation is always guaranteed to exist by Proposition 1(ii). By Theorem 2, under Assump. II.2, all of the $i$-th boundary equilibrium points are contained in the set $\mathcal{S}_{i}$ (defined in (33)) where the CLF and only the $i$-th CBF constraint are active in the QP (4). If the trajectory of the closed-loop system (3) is in the region of attraction of an asymptotically stable $i$-th boundary equilibrium point, then the state eventually enters $\mathcal{S}_{i}$. Then, the following strategy is considered:
(i) if the state is inside $\mathcal{S}_{i}$, $\overline{V}(x,\pi)$ must converge to $\overline{V}_{i}$, the closest $i$-th compatible CLF to $\overline{V}_{r}$, since this will induce a bifurcation on the closed-loop system state space, either removing or rendering the $i$-th boundary equilibrium points unstable.
(ii) if the state is outside $\cup^{N}_{i}\mathcal{S}_{i}$, $\overline{V}(x,\pi)$ must converge to the reference CLF $\overline{V}_{r}$. Therefore, if the state is in a deadlock-free region, its dynamics is determined from the reference CLF.
This desired effect can be achieved by the following QP controller for the CLF shape state $\pi\in\mathbb{R}^{\dim{\mathcal{S}^{n}}}$:
$u_{\pi}^{\star}$ $=\operatorname*{argmin}_{(u_{v},\delta_{v})\in\mathbb{R}^{p+1}}\lVert u_{\pi}\rVert^{2}+p_{\pi}\delta_{v}^{2}$ (73)
if $x\in\mathcal{S}_{i}$:
$\nabla V_{\pi}(\pi,H_{\overline{V}_{i}})^{\mathsf{T}}u_{\pi}+\gamma_{\pi}V_{\pi}(\pi,H_{\overline{V}_{i}})\leq\delta_{v}$
otherwise:
$\nabla V_{\pi}(\pi,H_{\overline{V}_{r}})^{\mathsf{T}}u_{\pi}+\gamma_{\pi}V_{\pi}(\pi,H_{\overline{V}_{r}})\leq\delta_{v}$
where $p_{\pi},\gamma_{\pi}>0$, set $\mathcal{S}_{i}$ as defined in (33) and $V_{\pi}$ as in (72) with reference CLFs drawn from the set $\{H_{\overline{V}_{r}},H_{\overline{V}_{1}},\cdots,H_{\overline{V}_{N}}\}$, depending on the region of the state space where the closed-loop state is located. With (73), the CLF shape state $\pi$ is controlled to achieve $\overline{V}\rightarrow\overline{V}_{i}$ when $x\in\mathcal{S}_{i}$, and $\overline{V}\rightarrow\overline{V}_{r}$ otherwise.
Remark V.3.
The controller proposed in (73) effectively controls the curvature of the CLF level sets in order to achieve CLF compatibility with respect to the $i$-th active barrier.
Remark V.4.
While the described strategy guarantees that stable equilibrium points are avoided, the occurrence of other types of attractors such as limit cycles is not theoretically eliminated.
V-C Numerical Simulation
In this section we present numerical examples demonstrating the viability of the proposed method. The code repository used for producing the results of this section is publicly available at https://github.com/CaipirUltron/CompatibleCLFCBF/tree/mydevel.
Consider again the two-dimensional LTI system $\dot{x}_{k}=-2x_{k}+u_{k}$, $k=1,2$, whose Q-function was shown in Fig. 1 for a given quadratic CLF $\overline{V}_{r}$ and CBF. This system satisfies Assump. III.1, and therefore no interior equilibrium points other than the origin exist.
Figure 2: CLF-CBF controller: fixed CLF vs adaptive strategy.
Here, we consider three quadratic barriers $h_{1}$, $h_{2}$ and $h_{3}$, with unsafe sets shown in Fig. 2 as red ellipses (left, top and right, respectively). The first row of Fig. 2 shows the results obtained using the nominal QP controller (4) with a fixed reference CLF $\overline{V}_{r}$ (level set is the blue ellipse) and all three CBF constraints. The Q-function shown in Fig. 1 was computed using CLF $\overline{V}_{r}$ and $h_{3}$, the CBF on the right. As expected, the trajectories converge towards a stable equilibrium point at $\partial\mathcal{C}_{3}$ (red dot). Two other unstable boundary equilibria also exist at $\partial\mathcal{C}_{3}$ (blue dots). Each of the remaining CBFs only have one unstable boundary equilibrium. Therefore, $\overline{V}_{r}$ is compatible with $h_{1}$ and $h_{2}$, and non-compatible with $h_{3}$.
Three compatible CLFs are computed using the optimization (70): $H_{\overline{V}_{1}}$, $H_{\overline{V}_{2}}$ and $H_{\overline{V}_{3}}$, each being the $i$-th compatible Hessian closest to the reference $H_{\overline{V}_{r}}$. Here, since $\overline{V}_{r}$ is already compatible with $h_{1}$ and $h_{2}$ (only unstable equilibrium points exist), $H_{\overline{V}_{1}}=H_{\overline{V}_{r}}$ and $H_{\overline{V}_{2}}=H_{\overline{V}_{r}}$. However, $H_{\overline{V}_{3}}$ is the Hessian of a CLF $\overline{V}_{3}$ (compatible with $h_{3}$) whose level sets are ellipses with slightly smaller eccentricity when compared to $\overline{V}_{r}$.
The second row of Fig. 2 shows the results obtained using our proposed compatible QP controller (4) with a CLF $V$ obtained from the inverse transformation of 7 of a quadratic CLF $\overline{V}(x,\pi)$, parametrized according to (71). In this example, we use the simple class $\mathcal{K}$ function $\gamma(V)=\gamma_{c}V$, where $\gamma_{c}>0$ is a constant. Solving (7), the transformed CLF is $\overline{V}(x,\pi)=\frac{1}{2}V^{2}$, and the inverse transformation is $V=2\overline{V}(x,\pi)^{\frac{1}{2}}$. The Hessian $H_{\overline{V}}(\pi)$ is controlled by our proposed strategy using (73). From the timestamps, the level sets of $\overline{V}$ dynamically change to match those of $\overline{V}_{3}$ when $x(t)\in\mathcal{S}_{3}$, inducing a bifurcation that removes the stable point (and one of the unstable points as well). Only one stable point remains at the boundary $\partial\mathcal{C}_{3}$. The system trajectories converge towards the origin for all tested initial conditions, and the level sets of $\overline{V}$ converge back to match those of $\overline{V}_{r}$ after the state leaves $\mathcal{S}_{3}$, as seen from the second row, third column of Fig. 2.
Figure 3: System trajectories using the proposed compatible CLF-CBF controller.
Now, consider another reference CLF $\overline{V}_{r}$, whose elliptical level set is shown in the first row of Fig. 3, with major-axis in the $y$-direction. In this case, $\overline{V}_{r}$ is not compatible with $h_{2}$ (the CBF on the top). Considering the nominal controller (4) with $\overline{V}_{r}$, the state converges towards the red stable point at the boundary $\partial\mathcal{C}_{2}$. However, once again using our proposed strategy, three compatible Hessians are computed from (70) and used in controller (73) to adapt the level sets of $\overline{V}(x,\pi)$ while the state is inside $\mathcal{S}_{2}$, effectively removing the stable equilibrium point from the second boundary. Once again, for all tested initial conditions, the system trajectories converge towards the origin and the CLF $\overline{V}$ converges towards $\overline{V}_{r}$ after the state leaves $\mathcal{S}_{2}$.
VI Conclusion
In this work, we have fully characterized the conditions for existence of undesirable equilibrium points arising in the CLF-CBF QP framework and their stability properties, considering affine nonlinear systems and multiple safety objectives. In particular, we have shown that the conditions for existence and instability of boundary equilibria depend on (10) and its derivatives (11). We demonstrate that boundary equilibrium points are always present for certain types of systems, and through the concept of CLF compatibility, we show that it is possible to choose the CLF in such a way that all stable boundary equilibrium points are removed. For driftless full-rank and LTI systems, the stability of the boundary equilibrium points can be studied using results from the theory of matrix polynomials, as described in Section V. Additionally, for this class of systems, we propose an algorithm for computing a compatible quadratic CLF with respect to a quadratic CBF, and a control strategy to modify the CLF geometry in (4), aiming to remove all stable equilibrium points from the closed-loop system.